Priya has spent $50,000 restoring her car and she plans to sell it for $200,000. However, in order for the car to be worth $200,000, she needs to spend another $60,000 to replace the engine. Without replacing the engine, she could sell the car at $110,000. Should she replace the engine before selling? Why, or why not? Explain your answer with reference to a principle of economics.
Answer #1

Cost for selling the car at $200000= 50000+60000 = 110000, the profits= 200000-110000 = 90000

If engine is not replaced, Profits = 110000-50000 = 60000

Therefore it is better for her to replace the engine to sell it for 200000 and earn a profit/revenue of 90000 which is 30000 more than the revenue when Priya doesn't replace the engine.